a) Calculate the amount of Calcium oxide required when it reacts with 852 gm of P4O10. 


 b) Write the structure of P4O10. 


Grade:11

1 Answers

Ruchi Shaw
10 years ago
6CaO + P4O10 ? 2Ca3(PO4)2
Moles of P4O10 = 852/284(Molecular mass) = 3
Moles of CaO = 3 × 6 = 18
Wt. of CaO = 18 × 56 = 1008 gm.

Structure of P4O10 :

Thanks & Regards :
Ruchi Shaw
askIITians faculty
MSc,Panjab University